Construct and explain a histogram by choosing an external law enforcement related data series. Examples of external data are reports published by the FBI or NIJ. include the following three elements:

  1. frequency distribution table for the data selected (includes cumulative frequency, relative frequency, and cumulative relative frequency)
  2. histogram graph based on the frequency distribution
  3. summary paragraph discussing observations and conclusions
Answer #1

Consider the following dataset showing the net increase in the external law enforcement personnel, for every 4-year period, from 1992-2012:

Years

Net increase(in thousands)

1992-1996 47
1996-2000 51
2000-2004 58
2004-2008 36
2008-2012 33

Note: This dataset is not an original dataset and is made up only for illustrative purposes. However, related original datasets can easliy be found.

Now, the frequency distribution for the given data is given below:

Frequency distribution
Years

Frequency (in thousands)

Relative Frequency

Cumulative Frequency

Cumulative Relative Frequecy

1992-1996 47 47/225 = 0.2089 47 0.2089
1996-2000 51 51/225 = 0.2267 47+51 = 98 0.2089+0.2267 = 0.4356
2000-2004 58 58/225 = 0.2578 98+58 = 156 0.4356+0.2578 = 0.6934
2004-2008 36 36/225 = 0.16 156+36 = 192 0.6934+0.16 = 0.8534
2008-2012 33 33/225 = 0.1466 192+33 = 225 0.8534+0.1466 = 1
Total 225 1

The calculations for relative frequencies, cumulative frequencies and cumulative relative frequencies are shown in the table itself.

The required histogram is:

Histogram showing net increase in enforcement personnel 70 60 50 40 30 20 1992-1996 1996-2000 2000-2004 2004-2008 2008-2012 Year

Observations & Conclusions:

A histogram is defined as the graphical representation of any numerical data in order to provide a better visual understanding of the distribution of data. On the other hand, a frequency distribution is the tabular representation of data.

The frequency distribution given above shows that there has been a total increase of 225,000 in the law enforcement personnel between 1992 - 2012. From the histogram, it can be clearly seen that maximum increase in the enforcement personnel happened during the period 200-2004, which is 58,000.It also shows that there has been a continuos increase from 1992 - 2004 but it showed a decline from 2004 - 2012. There is a significant decline in the period 2004 - 2008.